OpenStudy (anonymous):

find the values of sinpi/12 , cospi/12 , and tanpi/12 using the half angle formula.

OpenStudy (anonymous):

pi/4-pi/6=pi/12 so take the sin (pi/4-pi/6)

OpenStudy (anonymous):

sin(pi/4-pi/6)= sinpi/4cospi/6-cospi./4sinpi/6

OpenStudy (anonymous):